OpenStudy (anonymous):

expand 6 (big E lookin symbol) n=3 (5n+2)

OpenStudy (anonymous):

Is it this? \[\sum_{n=3}^{6} (5n+2)\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Ok, the letter is called Sigma, it's a greek letter. It tells us that we need to plug in the numbers from 3 to 6 (the number below the sigma and the number above the sigma) into the expression (5n+2). So we'll take 5n+2 and make n=3, n=4, n=5, and n=6, and then add them together. So I'll start you off with the first one, n=3: (5n+2)=(5(3) + 2)=15+2=17